OpenStudy (anonymous):

Can someone help me set up the equation(s) for this word problem please? A rectangular bin with an open top and volume of 38.72 cubic feet is to be built. The length of its base must be twice the width and the bin must be at least 3 feet high. Material for the base of the bin costs $12 per square foot and material for the sides costs $8 per square foot. If it costs $538.56 to build the bin, what are its dimensions?

OpenStudy (anonymous):

yikes so many words

OpenStudy (anonymous):

That's exactly my problem.

OpenStudy (anonymous):

lets call the width \(x\) and that makes the length \(2x\) the height is some unknown number \(y\) but it must be at least \(3\) the volume is length times width times height, which will now be \[2x^2y\]

OpenStudy (anonymous):

so one equation we can come up with is \[2x^2y=38.72\]

OpenStudy (anonymous):

the base has area \(2x^2\) at a cost of \(\$12\) per square foot, so it cost \(24x^2\)

OpenStudy (anonymous):

jeez this is a pain, but we are almost there

OpenStudy (anonymous):

there are four sides two have area \(xy\) and two have area \(2xy\) a picture would help, but i can't really draw one

OpenStudy (anonymous):



OpenStudy (anonymous):

wow you're a lot faster drawer than I am. XD

OpenStudy (anonymous):

so we get a cost of the sides of \[8xy+8xy+8\times 2xy+8\times 2xy\] \[=48xy\]

OpenStudy (anonymous):

for a total cost of \[24x^2+48xy\] which we know is \(538.56\) so we can set two equations \[24x^2+48xy=538.56\] and \[2x^2y=38.72\]

OpenStudy (anonymous):

i sincerely hope a) i didn't screw this up and b) it is making some sense to you before we go any further, any ideas?

OpenStudy (anonymous):

ooh i bet it is right, because this system has a nice solution !!

OpenStudy (anonymous):

So in summary the dimensions are 2x^2(y) = 38.72 where y >/= 3 and the price is 24x^2+ 48xy = 538.56?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now to solve ...

OpenStudy (anonymous):

you good with that or you need help solving it ?

OpenStudy (anonymous):

Could you tell me at least where to start with the dimensions? I don't really know what to do first.

OpenStudy (anonymous):

\[2x^2y=38.72\\ x^2y=19.36\\ y=\frac{19.36}{x^2}\]

OpenStudy (anonymous):

that solves the first equation for \(y\) now you can substitute in to the second equation

OpenStudy (anonymous):

24x^2+ 48xy = 538.56 24x^2+ 48x(19.36/x^2) = 538.56

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

and using that, you solve for x?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[24x^2+\frac{929.28}{x}=538.56\] is what i get

OpenStudy (anonymous):

although to be honest i just cheated at this point and asked for the answer to the system

OpenStudy (anonymous):

24x^2 + 48x(19.36/x^2) = 538.56 24x^2 + 929.28x / 48x^3 = 538.56 Get rid of an x x (24x + 929.28 / 48x^2) = 538.56 24x + 929.28 / 48x^2 = 538.56

OpenStudy (anonymous):

I'm confused O.o How did I get my x's so messed up?

OpenStudy (anonymous):

24x^2 + 929.28x / 48x^2 = 538.56 instead of 24x^2 + 929.28x / 48x^3 = 538.56 think

OpenStudy (anonymous):

then cancel an \(x\)

OpenStudy (anonymous):

Oh you only cancel an x from the fraction instead from the whole side of the equation.

OpenStudy (anonymous):

yeah i got to \[24x^2+\frac{929.28}{x}=538.56\]

OpenStudy (anonymous):

now i am not sure because we have to solve a cubic equation \[24x^2-538.56x+929.28=0\]

OpenStudy (anonymous):

no a cubic equation \[24x^3-538.56x+929.28=0\]

OpenStudy (anonymous):

not sure what method you are supposed to use like i said i just cheated

OpenStudy (anonymous):

on solution is apparently \(x=2.2\) which is the one you want because it makes \(y=4\)

OpenStudy (anonymous):

but i just cut to the chase at the beginning http://www.wolframalpha.com/input/?i=24x^2%2B48xy%3D538.56%2C2x^2y%3D38.72

OpenStudy (anonymous):

Glad to see I'm not the only one using wolfram.

OpenStudy (anonymous):

yeah i can't solve a cubic

OpenStudy (anonymous):

So of the solutions, how do I know which is correct?

OpenStudy (anonymous):

Do I have to plug each into the equation until I find the right one?

OpenStudy (anonymous):

the one that has \(y=4\) since you were told that \(y>3\)

OpenStudy (anonymous):

look at the second link

OpenStudy (anonymous):

Oh yeah. Duh. Forgot about that.

OpenStudy (anonymous):

Thank you so much for your help. With all those words and equations, I would have never been able to get the answer on my own. :)

OpenStudy (anonymous):

yw glad it worked out

OpenStudy (anonymous):

So in the end, the dimensions would be w= 2.2 l= 4.4 h= 4